# The tensor product of $p$-adic Hilbert spaces

### v1

@@ 4.1 Anti-unitary operators
Since $Z=J_{\Psi}U$ is the composition of the anti-isometric operator $J_{\Psi}$, and the unitary operator $U$, we observe that
@@
The implication (Z3) $\implies$ (Z4) is clear since, if $Z$ bounded, adjointable, and satisfies condition (90), then $Z$ is surjective and all-over. Moreover, if $\Psi\equiv\{\psi_{n}\}_{n\in\mathbb{N}}$ is an orthonormal basis in $\mathcal{H}$, by observing that
@@
for all $\chi,\eta$ in $\mathcal{H}$, we see that $Z$ is IP-conjugating too and, hence, (Z4) is verified. Next, let us prove (Z4) $\implies$ (Z5). Expressing $Z$ as $Z=J_{\Psi}L$ — for a suitable linear operator $L$ in $\mathcal{H}$, and with $J_{\Psi}$ the conjugation operator associated with the orthonormal basis $\Psi\equiv\{\psi_{n}\}_{n\in\mathbb{N}}$ — since, by assumption, $Z$ is surjective and IP-conjugating, we see that $L$ must be surjective and IP-preserving. Then, Theorem $5.8$ in [22] entails that $L$ is an adjointable top-linear isomorphism such that $L^{\ast}=L^{-1}$ and $\|L\|=\|L^{\ast}\|=\|L^{-1}\|$. This then yields
@@
Now, if $Z$ satisfies (Z5), expressing again $Z$ as $Z=J_{\Psi}L$, for some linear operator $L$, we see that, by hypothesis, $L$ must be an IP-preserving top-linear isomorphism such that $\|L\|=\|L^{-1}\|$. Then, by Lemma $5.2$ in [22], $L$ is an IP-reserving surjective isometry. Therefore, $Z=J_{\Psi}L$ is an IP-conjugating surjective anti-isometry, i.e., (Z5) $\implies$ (Z6). Let us now show (Z6) $\iff$ (Z7). Indeed, if $Z=J_{\Psi}L$ is anti-isometric, then $L$ is an isometry, and, hence, by Theorem $5.11$ in [22], an NO-preserving all-over operator. Taking into account that $J_{\Psi}$ is an anti-isometry, we see that $Z$ is an IP-conjugating, surjective, NO-preserving, all-over operator, i.e., (Z7) holds. Conversely, assume that $Z=J_{\Psi}L$ is IP-conjugating, surjective and all-over. Then, $L$ is a surjective IP-preserving operator and, hence, $\|L\|=\|L^{-1}\|$ (see Theorem $5.8$ in [22]). From this, it follows that
@@
As is well known, in a complex Hilbert space $\mathscr{K}$, every involutive, anti-unitary operator acts as a conjugation operator w.r.t. some orthonormal basis in $\mathscr{K}$ [28]. Specifically, if $Z$ is anti-unitary and satisfies $Z^{2}=\mathrm{Id}$, then an orthonormal basis $\Xi\equiv\{\xi_{k}\}_{k\in\mathbb{N}}$ in $\mathscr{K}$ exists such that $Z\xi_{k}=\xi_{k}$, for all $k\in\mathbb{N}$ — i.e., $Z\equiv J_{\Xi}$ is the conjugation operator associated with $\Xi\equiv\{\xi_{k}\}_{k\in\mathbb{N}}$ (see also Remark 4.11). In the next example, we show that, given an involutive anti-unitary operator $Z$ in a $p$-adic Hilbert space $\mathcal{H}$, one cannot always come up with an orthonormal basis $\Psi\equiv\{\psi_{m}\}_{m\in\mathbb{N}}$ in $\mathcal{H}$ preserved by $Z$.
@@
where $z_{1},z_{2}$ are suitably chosen coefficients in $\mathbb{Q}_{p}$. Note that, since $Z$ is anti-unitary, $\{Z\phi_{m}\}_{m\in\mathbb{N}}$ is an orthonormal set in $\mathcal{H}$ (see (Z8) in Theorem 4.8). By construction, $\Psi\equiv\{\psi_{m}\}_{m\in\mathbb{N}}$ is a $Z$-invariant dense subset of $\mathcal{H}$ — just note that $\phi_{m}=\frac{1}{2}z_{1}\psi_{2m}-\frac{\sqrt{\mu}}{2}z_{2}\psi_{2m-1}$. By suitably choosing the coefficients $z_{1},z_{2}$ in $\mathbb{Q}_{p}$, we can actually turn $\Psi\equiv\{\psi_{m}\}_{m\in\mathbb{N}}$ into a normal set. Specifically, let $z_{1}$ and $z_{2}$ be fixed in such a way that
$\max\big\{\left|1-z_{1}\right|,\left|z_{1}\right|\big\}\equiv t_{1}<1,\;\;\mbox{and}\;\;\max\bigg\{\left|1+\frac{z_{2}}{\sqrt{\mu}}\right|,\left|\frac{z_{2}}{\sqrt{\mu}}\right|\bigg\}\equiv t_{2}<1.$ (102)
With this choice of the coefficients $z_{1}$ and $z_{2}$, it is clear that $\Psi\equiv\{\psi_{m}\}_{m\in\mathbb{N}}$ is a normal basis in $\mathcal{H}$. Indeed, it suffices to notice that, for $m\in\mathbb{N}$ even, we have:
$\|\phi_{m}-\psi_{m}\|=\|\phi_{m}(1-z_{1})-z_{1}Z\phi_{m}\|\leq\max\big\{\left|1-z_{1}\right|,\left|z_{1}\right|\big\}\|\phi_{m}\|=t_{1}\|\phi_{m}\|<\|\phi_{m}\|=1,$ (103)
where, in the first inequality, we have used the fact that since $Z$ is an anti-unitary operator, it is isometric, i.e., $\|Z\phi_{m}\|=\|\phi_{m}\|$. Similarly, for $m\in\mathbb{N}$ odd, we find
$\|\phi_{m}-\psi_{m}\|=\bigg\|\bigg(1+\frac{z_{2}}{\sqrt{\mu}}\bigg)\phi_{m}-\frac{z_{2}}{\sqrt{\mu}}Z\phi_{m}\bigg\|\leq\max\bigg\{\left|1+\frac{z_{2}}{\sqrt{\mu}}\right|,\left|\frac{z_{2}}{\sqrt{\mu}}\right|\bigg\}\|\phi_{m}\|=t_{2}\|\phi_{m}\|<1.$ (104)
Therefore, setting $t\equiv\max\{t_{1},t_{2}\}$, from (103) and (104) and using the well known Full Perturbation Theorem for $t$-orthogonal bases in a $p$-adic normed space (see Theorem $2.3.16$ in [29]), we can argue that $\Psi\equiv\{\psi_{m}\}_{m\in\mathbb{N}}$ is a normal basis in $\mathcal{H}$. Next, let us see the constraints that $z_{1},z_{2}$ have to satisfy for $\Psi\equiv\{\psi_{m}\}_{m\in\mathbb{N}}$ to be IP-orthonormal too. It is clear that $\langle\psi_{n},\psi_{m}\rangle=0$, for all $n\neq m\in\mathbb{N}$, i.e., $\Psi$ is IP-orthogonal; requiring further the IP-orthonormality we get the following conditions:
$1=\langle\psi_{m},\psi_{m}\rangle=\langle z_{1}(Z\phi_{m}+\phi_{m}),z_{1}(Z\phi_{m}+\phi_{m})\rangle=2z_{1}^{2}$ (105)
@@
i.e., $z_{1}$, $z_{2}$ do not belong to $\mathbb{Q}_{p}$. On the other hand, if, in the defining condition of the set $\Psi\equiv\{\psi_{m}\}_{m\in\mathbb{N}}$ in (101) above, we assume that $z_{1},z_{2}$ are in $\mathbb{Q}_{p,\mu}$ — i.e., $z_{1}$ and $z_{2}$ belong to a suitable quadratic extension of $\mathbb{Q}_{p}$ — setting $z_{1}$ and $z_{2}$ as in (107), we can turn $\Psi\equiv\{\psi_{m}\}_{m\in\mathbb{N}}$ into an orthonormal basis; but, in this case, it will not be an invariant set under the action of $Z$. Concluding, we have the following dichotomy: Either $\Psi\equiv\{\psi_{m}\}_{m\in\mathbb{N}}$ is a $Z$-invariant, normal, IP-orthogonal basis — in which case it is not IP-orthonormal — or $\Psi\equiv\{\psi_{m}\}_{m\in\mathbb{N}}$ is an orthonormal basis, but not $Z$-invariant.
In example 4.9 we construct an invariant dense subset of $\mathcal{H}$, which cannot be turned into an IP-orthonormal set. Actually, this example is a manifestation of a peculiar feature of the conjugation operators in the $p$-adic setting, as we will see in the next Proposition 4.10. Before stating this result, let us first note the following. Consider the canonical Hilbert space $\mathbb{H}$ endowed with the (so-called) canonical basis $\bm{e}\equiv\{e_{i}\}_{i\in\mathbb{N}}$ (see Example 2.11). Define in $\mathbb{H}$ the transformation $J_{0}$ by setting
@@
where $\mathcal{P}\colon\mathbb{H}\rightarrow c_{0}(\mathbb{N},\mathbb{Q}_{p,\mu})$ is the map defined as
@@
i.e., $J_{\Psi}\psi_{n}=W_{\Psi}^{-1}J_{0}W_{\Psi}\psi_{n}=\psi_{n}$. We have thus proved that, given any orthonormal basis $\Psi\equiv\{\psi_{n}\}_{n\in\mathbb{N}}$ in $\mathcal{H}$, it is always possible to construct a conjugation operator associated with it; moreover, the so-constructed conjugation operator is equivalent, up to isomorphisms, with the conjugation $J_{0}$ in the canonical Hilbert space $\mathbb{H}$. A natural question to ask is whether the reverse implication holds, that is, if, given an involutive anti-unitary operator $Z$ in $\mathcal{H}$, there always exists an Hilbert space isomorphism mapping $Z$ into the canonical conjugation operator $J_{0}$. The next result shows that, in fact, it is not so.
@@
with $Z\chi_{1}=\chi_{1}$ and $Z\chi_{2}=\chi_{2}$, i.e., $\chi_{1},\chi_{2}$ are $Z$-invariant. Thus, every $\chi$ in $\mathcal{H}$ can be expressed as a linear combination of elements in $\mathfrak{S}$. It then follows that a $Z$-invariant orthonormal basis in $\mathcal{H}$ exists if and only if we can construct an orthonormal basis in $\mathfrak{S}$. Let $\Phi\equiv\{\phi_{n}\}_{n\in\mathbb{N}}$ be an orthonormal basis in $\mathcal{H}$. It is clear that a $Z$-invariant dense subset $\Psi\equiv\{\psi_{n}\}_{n\in\mathbb{N}}$ in $\mathcal{H}$ — i.e., a dense subset in $\mathfrak{S}$ — can be constructed by setting
@@
A conjugation operator $J$, in a complex Hilbert space $\mathscr{K}$, is defined as an involutive anti-unitary operator. Then, it is not difficult to prove that there exists an orthonormal basis $\Psi\equiv\{\psi_{k}\}_{k\in\mathbb{N}}$ in $\mathscr{K}$ which is invariant under the action of $J$, i.e., such that $J\psi_{k}=\psi_{k}$, for all $k\in\mathbb{N}$ [28]. Indeed, let $\mathfrak{G}$ denote the set of all vectors $\chi$ in $\mathscr{K}$ invariant under the action of $J$, i.e., such that $J\chi=\chi$; this set is easily seen to be dense in $\mathscr{K}$. Since it is non-empty, it is separable and, thus, contains an everywhere dense subset. Applying to this set the so-called Gram–Schmidt orthonormalization process yields a complete orthonormal subset of $\mathfrak{G}$ and — since $\mathfrak{G}$ is dense — a $J$-invariant basis in $\mathscr{K}$. In the $p$-adic setting, what prevents us to repeat a similar general construction is the lack of a $p$-adic counterpart of the Gram–Schmidt process, which precludes the possibility to turn a normal set into an orthonormal basis in the invariant subspace — thus, an invariant orthonormal basis in the Hilbert space.
@@ 4.2 Construction of the isomorphism
Let $\Phi=\{\phi_{i}\}_{i\in\mathbb{N}}$ and $\Psi=\{\psi_{j}\}_{j\in\mathbb{N}}$ be two orthonormal bases in the $p$-adic Hilbert spaces $\mathcal{H}$ and $\mathcal{K}$ respectively. Remember that $\{{{}^{ji}}E^{\Phi,\Psi}\}_{j,i\in\mathbb{N}}=\{\lvert\psi_{j}\rangle\langle\phi_{i}\rvert\}_{j,i\in\mathbb{N}}\equiv\{\langle\phi_{i},\cdot\rangle\psi_{j}\}_{j,i\in\mathbb{N}}$ is an orthonormal basis for $\mathcal{T}(\mathcal{H},\mathcal{K})$ (see Theorem 2.24). Therefore, any $T\in\mathcal{T}(\mathcal{H},\mathcal{K})$ can be written, in Dirac notation, as
$T=\sum_{j,i\in\mathbb{N}}T_{ji}\lvert\psi_{j}\rangle\langle\phi_{i}\rvert,\quad(T_{ji})\in\mathsf{M}_{\infty}(\mathbb{Q}_{p,\mu}).$ (118)
For an operator $T\in\mathcal{T}(\mathcal{H},\mathcal{K})$, we also have that
$\langle{{}^{ji}}E^{\Phi,\Psi},T\rangle_{\mbox{\tiny$\mathcal{T}(\mathcal{H},\mathcal{K})$}}=\mathrm{tr}(|\psi_{j}\rangle\langle\phi_{i}|T)=T_{ji}.$ (119)
@@
Let $J_{\Phi}$ be the conjugation operator associated with the orthonormal basis $\Phi\equiv\{\phi_{i}\}_{i\in\mathbb{N}}\subset\mathcal{H}$ (cf. Definition 4.2), and let $\Psi\equiv\{\psi_{j}\}_{j\in\mathbb{N}}$ be an orthonormal basis in $\mathcal{K}$. Let us define the map
@@
$\mathcal{I}(\lvert w\rangle\langle v\rvert)=(J_{\Phi}v)\otimes w$ (121)
holds for every vector $v=\sum_{i\in\mathbb{N}}\langle\phi_{i},v\rangle\phi_{i}$ in $\mathcal{H}$ and $w=\sum_{j\in\mathbb{N}}\langle\psi_{j},v\rangle\psi_{j}$ in $\mathcal{K}$, such that $\lvert w\rangle\langle v\rvert\in\mathcal{T}(\mathcal{H},\mathcal{K})$.
@@
[…]
@@
Next result provides the desired isomorphism between the tensor product of $p$-adic Hilbert spaces $\mathcal{H}\otimes\mathcal{K}$, and the Hilbert space of $p$-adic Hilbert-Schmidt operators $\mathcal{T}(\mathcal{H},\mathcal{K})$.
@@
$\mathcal{I^{*}}\colon\mathcal{H}\otimes\mathcal{K}\ni u\mapsto\mathcal{I^{*}}(u)\coloneqq\sum_{i,j\in\mathbb{N}}\langle\phi_{i}\otimes\psi_{j},u\rangle\lvert\psi_{j}\rangle\langle\phi_{i}\rvert\in\mathcal{T}(\mathcal{H},\mathcal{K})$ (123)
@@
Since $\mathcal{I}$ is linear and maps the norm-orthogonal basis $\{\lvert\psi_{j}\rangle\langle\phi_{i}\rvert\}_{j,i\in\mathbb{N}}$ to $\{(J_{\Phi}\phi_{i})\otimes\psi_{j}\}_{i,j\in\mathbb{N}},$ then it is inner-product preserving and injective. To prove it is also surjective, let us take $u\in\mathcal{H}\otimes\mathcal{K}$, and let $T\in\mathcal{T}(\mathcal{H},\mathcal{K})$ be written as
$T=\sum_{i,j\in\mathbb{N}}\langle\phi_{j}\otimes\psi_{i},u\rangle\lvert\psi_{i}\rangle\langle\phi_{j}\rvert$. (124)
@@
$\begin{split}\mathcal{I}(T)&\coloneqq\sum_{i,j\in\mathbb{N}}\langle\psi_{j},T\phi_{i}\rangle(\phi_{i}\otimes\psi_{j})=\sum_{i,j\in\mathbb{N}}\langle\psi_{j},\sum_{s,t\in\mathbb{N}}\langle\phi_{s}\otimes\psi_{t},u\rangle\psi_{t}\rangle\langle\phi_{s},\phi_{i}\rangle(\phi_{i}\otimes\psi_{j})\\ &\>=\sum_{s,t\in\mathbb{N}}\langle\phi_{s}\otimes\psi_{t},u\rangle\sum_{i,j\in\mathbb{N}}\delta_{j,t}\delta_{s,i}(\phi_{i}\otimes\psi_{j})=\sum_{i,j\in\mathbb{N}}\langle\phi_{i}\otimes\psi_{j},u\rangle(\phi_{i}\otimes\psi_{j})=u.\end{split}$ (125)
@@
$\bigg\langle T,\sum_{i,j\in\mathbb{N}}\langle\phi_{i}\otimes\psi_{j},u\rangle\lvert\psi_{j}\rangle\langle\phi_{i}\rvert\bigg\rangle_{\mbox{\tiny$\mathcal{T}(\mathcal{H},\mathcal{K})$}}=\langle\mathcal{I}(T),u\rangle=\langle T,\mathcal{I}^{*}(u)\rangle.$ (128)
@@
For the operator $\lvert w\rangle\langle v\rvert\in\mathcal{T}(\mathcal{H},\mathcal{K})$ — where $v\in\mathcal{H}$ and $w\in\mathcal{K}$ are any two vectors — the following equalities hold

### v2

@@ 4.1 Anti-unitary operators
Moreover, since $Z=J_{\Psi}U$, and $J_{\Psi}$ is an anti-isometric operator, we observe that
@@
The implication (Z3) $\implies$ (Z4) is clear since, if $Z$ is bounded, adjointable, and satisfies condition (92), then $Z$ is surjective and all-over. Moreover, if $\Psi\equiv\{\psi_{n}\}_{n\in\mathbb{N}}$ is an orthonormal basis in $\mathcal{H}$, by observing that
@@
for all $\chi,\eta$ in $\mathcal{H}$, we see that $Z$ is IP-conjugating too and, hence, (Z4) is verified. Next, let us prove (Z4) $\implies$ (Z5). Expressing $Z$ as $Z=J_{\Psi}L$ — for a suitable linear operator $L$ in $\mathcal{H}$, and with $J_{\Psi}$ the conjugation operator associated with the orthonormal basis $\Psi\equiv\{\psi_{n}\}_{n\in\mathbb{N}}$ — since, by assumption, $Z$ is surjective and IP-conjugating, we see that $L$ must be surjective and IP-preserving. Then, Theorem $5.8$ in [22] allows us to infer that $L$ is an adjointable top-linear isomorphism such that $L^{\ast}=L^{-1}$ and $\|L\|=\|L^{\ast}\|=\|L^{-1}\|$. Hence, we have:
@@
Now, if $Z$ satisfies (Z5), expressing again $Z$ as $Z=J_{\Psi}L$, for some linear operator $L$ on $\mathcal{H}$, we see that, by hypothesis, $L$ must be an IP-preserving top-linear isomorphism such that $\|L\|=\|L^{-1}\|$. Then, by Lemma $5.2$ in [22], $L$ is an IP-reserving surjective isometry. Therefore, $Z=J_{\Psi}L$ is an IP-conjugating surjective anti-isometry, i.e., (Z5) $\implies$ (Z6). Let us now show (Z6) $\iff$ (Z7). Indeed, if $Z=J_{\Psi}L$ is anti-isometric, then $L$ is an isometry, and, by Theorem $5.11$ in [22], a NO-preserving all-over operator. Recalling that $J_{\Psi}$ is an anti-isometry (see (i) in Proposition 4.6), we see that $Z$ is an IP-conjugating, surjective, NO-preserving, all-over operator, i.e., (Z7) holds. Conversely, assume that $Z=J_{\Psi}L$ is IP-conjugating, surjective and all-over. Then, $L$ is a surjective IP-preserving operator and, hence, $\|L\|=\|L^{-1}\|$ (cf. Theorem $5.8$ in [22]). From this, it follows that
@@
As is well known, in a complex Hilbert space $\mathscr{K}$, every involutive, anti-unitary operator acts as a conjugation operator w.r.t. some orthonormal basis in $\mathscr{K}$ [28]. Specifically, if $Z$ is anti-unitary and satisfies $Z^{2}=\mathrm{Id}$, then an orthonormal basis $\mathscr{Y}\equiv\{y_{k}\}_{k\in\mathbb{N}}$ in $\mathscr{K}$ exists such that $Zy_{k}=y_{k}$, for all $k\in\mathbb{N}$ — i.e., $Z\equiv J_{\mathscr{Y}}$ is the conjugation operator associated with $\mathscr{Y}\equiv\{y_{k}\}_{k\in\mathbb{N}}$ (see also Remark 4.11). In the next example, we show that this picture does not extend to the $p$-adic framework, namely, that given an involutive anti-unitary operator $Z$ in a $p$-adic Hilbert space $\mathcal{H}$, one cannot always come up with an orthonormal basis $\Psi\equiv\{\psi_{m}\}_{m\in\mathbb{N}}$ in $\mathcal{H}$ preserved by $Z$.
@@
where $z_{1},z_{2}$ are suitably chosen coefficients in $\mathbb{Q}_{p}$. Note that, since $Z$ is anti-unitary, $\{Z\phi_{m}\}_{m\in\mathbb{N}}$ is an orthonormal set in $\mathcal{H}$ (see (Z8) in Theorem 4.8). By construction, $\Psi\equiv\{\psi_{m}\}_{m\in\mathbb{N}}$ is a $Z$-invariant dense subset of $\mathcal{H}$ — just note that $\phi_{m}=\frac{1}{2}z_{1}\psi_{2m}-\frac{\sqrt{\mu}}{2}z_{2}\psi_{2m-1}$. Moreover, by suitably choosing the coefficients $z_{1},z_{2}$ in $\mathbb{Q}_{p}$, we can turn $\Psi$ into a normal subset of $\mathcal{H}$. Next, let us see the constraints that $z_{1},z_{2}$ have to satisfy for $\Psi\equiv\{\psi_{m}\}_{m\in\mathbb{N}}$ to be orthonormal too. Clearly, $\langle\psi_{n},\psi_{m}\rangle=0$, for all $n\neq m\in\mathbb{N}$, i.e., $\Psi$ is IP-orthogonal; requiring further the IP-orthonormality we get the following conditions:
$1=\langle\psi_{m},\psi_{m}\rangle=\langle z_{1}(Z\phi_{m}+\phi_{m}),z_{1}(Z\phi_{m}+\phi_{m})\rangle=2z_{1}^{2},$ (104)
@@
i.e., $z_{1}$, $z_{2}$ do not belong to $\mathbb{Q}_{p}$. On the other hand, if, in the defining condition of the set $\Psi\equiv\{\psi_{m}\}_{m\in\mathbb{N}}$ in (103) above, we assume that $z_{1},z_{2}$ are in $\mathbb{Q}_{p,\mu}$ — i.e., $z_{1}$ and $z_{2}$ belong to a suitable quadratic extension of $\mathbb{Q}_{p}$ — setting $z_{1}$ and $z_{2}$ as in (106), we can turn $\Psi\equiv\{\psi_{m}\}_{m\in\mathbb{N}}$ into an orthonormal subset of $\mathcal{H}$; but, in this case, it will not be an invariant set under the action of $Z$. Concluding, we have the following dichotomy: Either $\Psi\equiv\{\psi_{m}\}_{m\in\mathbb{N}}$ is a $Z$-invariant, normal, IP-orthogonal dense subset of $\mathcal{H}$ — in which case it is not IP-orthonormal — or $\Psi\equiv\{\psi_{m}\}_{m\in\mathbb{N}}$ is an orthonormal dense subset, but not $Z$-invariant.
In example 4.9, we constructed an invariant dense subset of $\mathcal{H}$, which cannot be turned into an orthonormal set. Actually, this example is a manifestation of a peculiar feature of the conjugation operators in the $p$-adic setting, as will be clarified in the forthcoming Proposition 4.10. Before stating this result, let us first note the following. Consider the canonical Hilbert space $\mathbb{H}$ endowed with the (so-called) canonical basis $\bm{e}\equiv\{e_{i}\}_{i\in\mathbb{N}}$ (see Example 2.11). Define in $\mathbb{H}$ the transformation $J_{0}$ by setting
@@
where $\mathcal{P}\colon\mathbb{H}\rightarrow c_{0}(\mathbb{N},\mathbb{Q}_{p,\mu})$ is the natural embedding of $\mathbb{H}$ into $c_{0}(\mathbb{N},\mathbb{Q}_{p,\mu})$ given by
@@
i.e., $J_{\Psi}\psi_{n}=W_{\Psi}^{-1}J_{0}W_{\Psi}\psi_{n}=\psi_{n}$. We have thus proved that, given any orthonormal basis $\Psi\equiv\{\psi_{n}\}_{n\in\mathbb{N}}$ in $\mathcal{H}$, it is always possible to construct a conjugation operator associated with it; moreover, the so-constructed conjugation operator is equivalent, up to isomorphisms, with the conjugation $J_{0}$ in the canonical Hilbert space $\mathbb{H}$. A natural question to ask is whether the reverse implication holds, namely if, given an involutive anti-unitary operator $Z$ in $\mathcal{H}$, there always exists an Hilbert space isomorphism $\mathscr{I}$ such that $\mathscr{I}^{-1}Z\mathscr{I}$ reduces to the canonical conjugation operator $J_{0}$. The next result shows that, in fact, it is not so.
@@
with $Z\chi_{1}=\chi_{1}$ and $Z\chi_{2}=\chi_{2}$, i.e., $\chi_{1},\chi_{2}\in\mathfrak{S}$. Thus, every $\chi$ in $\mathcal{H}$ can be expressed as a linear combination of elements in $\mathfrak{S}$. It then follows that a $Z$-invariant orthonormal basis in $\mathcal{H}$ exists if and only if we can construct an orthonormal basis in $\mathfrak{S}$. Let $\Phi\equiv\{\phi_{n}\}_{n\in\mathbb{N}}$ be an orthonormal basis in $\mathcal{H}$. A $Z$-invariant dense subset $\Psi\equiv\{\psi_{n}\}_{n\in\mathbb{N}}$ in $\mathcal{H}$ — i.e., a dense subset in $\mathfrak{S}$ — can be constructed by setting
@@
A conjugation operator $J$ in a complex Hilbert space $\mathscr{K}$ is defined as an involutive anti-unitary operator. Then, it is not difficult to prove that there exists an orthonormal basis $\Psi\equiv\{\psi_{k}\}_{k\in\mathbb{N}}$ in $\mathscr{K}$ which is invariant under the action of $J$, i.e., such that $J\psi_{k}=\psi_{k}$, for all $k\in\mathbb{N}$ [28]. Indeed, let $\mathfrak{G}$ denote the set of all vectors $\chi$ in $\mathscr{K}$ invariant under the action of $J$, i.e., such that $J\chi=\chi$; this set is easily seen to be dense in $\mathscr{K}$. Since it is non-empty, it is separable and, thus, contains an everywhere dense subset (see Theorem $1.18$ in [36]). Applying to this set the so-called Gram–Schmidt orthonormalization process yields a complete orthonormal subset of $\mathfrak{G}$ and — since $\mathfrak{G}$ is dense — a $J$-invariant basis in $\mathscr{K}$. In the $p$-adic setting, what prevents us to repeat a similar general construction is the lack of a $p$-adic counterpart of the Gram–Schmidt process, which precludes the possibility to turn a normal set into an orthonormal basis in the invariant subspace — thus, an invariant orthonormal basis in the Hilbert space.
@@ 4.2 Construction of the isomorphism
We now proceed to construct an isomorphism between the $p$-adic tensor product Hilbert space $\mathcal{H}\otimes\mathcal{K}$ and the space $\mathcal{T}(\mathcal{H},\mathcal{K})$ of the trace/Hilbert-Schmidt class operators. This is not simply an ‘additional result’ but, rather, a consistency check that our previous construction of the tensor product of two $p$-adic Hilbert spaces — as presented in Section 3 — provides indeed the correct counterpart to the usual tensor product of the complex setting.
Thus, let $\Phi=\{\phi_{i}\}_{i\in\mathbb{N}}$ and $\Psi=\{\psi_{j}\}_{j\in\mathbb{N}}$ be two orthonormal bases in the $p$-adic Hilbert spaces $\mathcal{H}$ and $\mathcal{K}$, respectively. Recall that $\{{{}^{ji}}E^{\Phi,\Psi}\}_{j,i\in\mathbb{N}}=\{\lvert\psi_{j}\rangle\langle\phi_{i}\rvert\}_{j,i\in\mathbb{N}}\equiv\{\langle\phi_{i},\cdot\rangle\psi_{j}\}_{j,i\in\mathbb{N}}$ is an orthonormal basis in $\mathcal{T}(\mathcal{H},\mathcal{K})$ (see Theorem 2.24); hence, any $T\in\mathcal{T}(\mathcal{H},\mathcal{K})$ can be written, in a unique way, as
$T=\sum_{j,i\in\mathbb{N}}T_{ji}\langle\phi_{i},\cdot\hskip 1.13809pt\rangle\psi_{j},\quad\mbox{where}\quad T_{ji}=\langle{{}^{ji}}E^{\Phi,\Psi},T\rangle_{\mbox{\tiny$\mathcal{T}(\mathcal{H},\mathcal{K})$}}=\mathrm{tr}(\langle\phi_{i},\cdot\hskip 1.13809pt\rangle\psi_{j}T).$ (117)
@@
Let $\mathcal{H}$ and $\mathcal{K}$ be two $p$-adic Hilbert spaces, let $J_{\Phi}$ be the conjugation operator associated with the orthonormal basis $\Phi\equiv\{\phi_{i}\}_{i\in\mathbb{N}}\subset\mathcal{H}$ (cf. Definition 4.2), and let $\Psi\equiv\{\psi_{j}\}_{j\in\mathbb{N}}$ be an orthonormal basis in $\mathcal{K}$. Define the map
@@
$\mathcal{I}(\langle v,\cdot\hskip 1.13809pt\rangle w)=(J_{\Phi}v)\otimes w$ (119)
holds for every vector $v=\sum_{i\in\mathbb{N}}\langle\phi_{i},v\rangle\phi_{i}$ in $\mathcal{H}$, and $w=\sum_{j\in\mathbb{N}}\langle\psi_{j},v\rangle\psi_{j}$ in $\mathcal{K}$, such that $\langle v,\cdot\hskip 1.13809pt\rangle w\in\mathcal{T}(\mathcal{H},\mathcal{K})$.
@@
$\mathcal{I}(\langle v,\cdot\hskip 1.13809pt\rangle w)$ $=\sum_{i,j\in\mathbb{N}}\langle\psi_{j},w\rangle\langle v,\phi_{i}\rangle(\phi_{i}\otimes\psi_{j})$
@@
Next result provides the desired isomorphism between the tensor product of $p$-adic Hilbert spaces $\mathcal{H}\otimes\mathcal{K}$, and $\mathcal{T}(\mathcal{H},\mathcal{K})$.
@@
$\mathcal{I^{*}}\colon\mathcal{H}\otimes\mathcal{K}\ni u\mapsto\mathcal{I^{*}}(u)\coloneqq\sum_{i,j\in\mathbb{N}}\langle\phi_{i}\otimes\psi_{j},u\rangle\langle\phi_{i},\cdot\hskip 1.13809pt\rangle\psi_{j}\in\mathcal{T}(\mathcal{H},\mathcal{K})$ (121)
@@
Since $\mathcal{I}$ is linear and maps the norm-orthogonal basis $\{\langle\phi_{i},\cdot\hskip 1.13809pt\rangle\psi_{j}\}_{j,i\in\mathbb{N}}$ to $\{(J_{\Phi}\phi_{i})\otimes\psi_{j}\}_{i,j\in\mathbb{N}},$ then it is IP-preserving and injective. To prove it is also surjective, let us take $u\in\mathcal{H}\otimes\mathcal{K}$, and let $T\in\mathcal{T}(\mathcal{H},\mathcal{K})$ be written as
$T=\sum_{i,j\in\mathbb{N}}\langle\phi_{j}\otimes\psi_{i},u\rangle\langle\phi_{j},\cdot\hskip 1.13809pt\rangle\psi_{i}.$ (122)
@@
$\begin{split}\mathcal{I}(T)&\coloneqq\sum_{i,j\in\mathbb{N}}\langle\psi_{j},T\phi_{i}\rangle(\phi_{i}\otimes\psi_{j})=\sum_{i,j\in\mathbb{N}}\bigg\langle\psi_{j},\sum_{s,t\in\mathbb{N}}\langle\phi_{s}\otimes\psi_{t},u\rangle\psi_{t}\bigg\rangle\langle\phi_{s},\phi_{i}\rangle(\phi_{i}\otimes\psi_{j})\\ &\>=\sum_{s,t\in\mathbb{N}}\langle\phi_{s}\otimes\psi_{t},u\rangle\sum_{i,j\in\mathbb{N}}\delta_{jt}\delta_{si}(\phi_{i}\otimes\psi_{j})=\sum_{i,j\in\mathbb{N}}\langle\phi_{i}\otimes\psi_{j},u\rangle(\phi_{i}\otimes\psi_{j})=u.\end{split}$ (123)
@@
$\bigg\langle T,\sum_{i,j\in\mathbb{N}}\langle\phi_{i}\otimes\psi_{j},u\rangle\langle\phi_{i},\cdot\hskip 1.13809pt\rangle\psi_{j}\bigg\rangle_{\mbox{\tiny$\mathcal{T}(\mathcal{H},\mathcal{K})$}}=\langle\mathcal{I}(T),u\rangle=\langle T,\mathcal{I}^{*}(u)\rangle.$ (126)
@@
To conclude this section we collect, in the following result, some useful consequences of Theorem 4.14
@@
Let $\mathcal{H}$ and $\mathcal{K}$ be two $p$-adic Hilbert spaces, and let $J_{\Phi}$ be the conjugation operator associated with the orthonormal basis $\Phi\equiv\{\phi_{n}\}_{n\in\mathbb{N}}\subset\mathcal{H}$. Then, for the operator $\langle v,\cdot\hskip 1.13809pt\rangle w\in\mathcal{T}(\mathcal{H},\mathcal{K})$ — where $v\in\mathcal{H}$ and $w\in\mathcal{K}$ are any two vectors — the following equalities hold